An aqueous solution containing 6.33 g of an optically pure compound was diluted to 500.0 mL with water and placed in a polarimeter tube 10.0 cm long. The measured rotation was –5.91° at 25 °C. Calculate the specific rotation of the compound.
Answer #1

The specific rotation of a compound is defined as:

[]D=

where $[\alpha]_{\mathrm{D}}$ is the specific rotation in degrees per millimeter per gram, $\alpha$ is the observed rotation in degrees, $c$ is the concentration in g/mL, and $l$ is the path length in decimeters.

We are given that the solution contains 6.33 g of the compound in 500.0 mL of solution, which corresponds to a concentration of:

=6.33g0.5000L=12.66g/L

We are also given that the measured rotation is –5.91° in a polarimeter tube with a length of 10.0 cm, or 0.1 dm. Thus, we can calculate the specific rotation as:

[]D=5.9112.66g/L0.1dm=46.6ml/g

Therefore, the specific rotation of the compound is -46.6 degrees per millimeter per gram.


answered by: Hydra Master